reecej123:

The side length, s, of a cube is x – 2y. If V = s3, what is the volume of the cube? x3 – 6x2y + 12xy2 – 8y3 x3 + 6x2y + 12xy2 + 8y3 3x3 – 6x2y + 12xy2 – 24y3 3x3 + 6x2y + 12xy2 + 24y3

jhonyy9:

so volume = s^3 what mean (x-2y)^3 do you know formula (a-b)^3 = ?

reecej123:

no

jhonyy9:

ok than how you think what mean (a-b)^3 ?

reecej123:

idk im really confused

jhonyy9:

ok what mean 2^3 ?

reecej123:

8

jhonyy9:

or what mean 2^2 ?

jhonyy9:

yes but how you ve got this 8 ?

reecej123:

2^3 is 8 which means 2^2 is 4

jhonyy9:

yes this is true but how you get these 8 and 4 ?

reecej123:

not quite sure

jhonyy9:

2^2 = 2*2 = 4 bc. the exponent is 2 so this mean that you need multiply 2 time the 2 what is 2*2 = 4 and this way with 2^3 multiply 2 by 2 ,3 time

reecej123:

confused again

TheSmartOne:

He's basically saying that 2^2 is 2*2 You multiply it by itself twice because the exponent is 2 When you do 2^3, since the exponent is 3, you multiply 2 by itself three times so 2*2*2 = 8 2^3 = 8 Does that make sense?

TheSmartOne:

Now how does this relate to your question? Your question involves a cube and they said that the side of the cube is (x-2y) To find the volume of this cube, we have to use the formula they gave which is V = s^3 Plug in what they gave us for s which is x-2y And you get V = (x-2y)^3 Now obviously we have to expand this and find out the correct answer.

TheSmartOne:

So there's two ways to go about answering it You don't know the formula for (a-b)^3 And I'll give that to you in a minute, would be handy to have it memorized The other method is the longer way which you manually multiply it out and you have to use all that distributing rules you learned. And this is actually how they came up with the formula (a-b)^3 = (a-b)(a-b)(a-b) = (a^2 - 2ab + b^2)(a-b) \(= \boxed{a^3 - 3a^2b + 3ab^2 - b^3}\) I skipped a few steps but like that's how the formula is derived. So like if you want you could do (x-2y)(x-2y)(x-2y) And you can multiply it all out Or you can use the formula and x would be the a, and 2y is your b Let me know if you have any questions or would like more clarification!